OpenStudy (anonymous):

Solve the inequality: 150 ≤ 30x + 10x ≤ 200

OpenStudy (anonymous):

30x+10x is the same as 40x. So 40x has to be greater than or equal to 150, but less than or equal to 200, so \[x \ge150/40=3.75, x \le200/40=5\] so\[3.75\le x \le5\]